OpenStudy (anonymous):

Lucy rode on a train that traveled at a constant rate. The graph compares the number of hours she traveled to the number of miles she traveled. What was the rate of speed of the train? A. 50 miles per hour B. 75 miles per hour C. 100 miles per hour D. 125 miles per hour

OpenStudy (imstuck):

Do you have an image of the graph to post for us?

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

Do you know how to ind the equation of this line?

OpenStudy (anonymous):

Just take two points on the graph and plug them into m = y2 - y1/x2 - x1 We can use (2, 150) and (4, 300) \[m = \frac{ y2 - y1 }{ x2 - x1 }\] \[m = \frac{ 300 - 150 }{ 4 - 2 }\] \[m = \frac{ 150 }{ 2 }\] \[m = 75\]

OpenStudy (anonymous):

So the train went 75 MPH.

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

i have more can u help??

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

The graph shows the amount of commission that Ms. Hayes received. Which describes the rate of change in the graph? A. As sales increase by $1,000, commissions increase by $100. B. As sales increase by $1,000, commissions increase by $150. C. As sales increase by $1,000, commissions increase by $200. D. The graph does not show a constant rate of change.

OpenStudy (anonymous):

It's obviously B.

OpenStudy (anonymous):

\[(1000, 150) (2000, 300) (3000, 450) (4000, 600) (5000, 750) (6000, 900)\] x increases by 1000 y increases by 150

OpenStudy (anonymous):

No chart for this one Which situation can be represented by the equation 3s − 25 = 80? A. Marisa bought 25 sweaters for $3 each with a discount of s for a bill of $80. B. Marisa bought 3 sweaters for s dollars each with a $25 discount for a bill of $80. C. Marisa bought s sweaters for $25 each with a $3 discount for a bill of $80. D. Marisa bought 3 sweaters for s dollars and paid $25 in tax for a bill of $80.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

A membership at a modern art museum costs $85 per year. Members pay a reduced price of $7.50 for entrance into each special exhibition. Which of the following statements is true of the annual cost of a membership at the museum? Let y = total cost and x = number of special exhibitions seen. A. The annual cost is shown by the linear function y = 85x + 7.50. The rate of change is 85, and the initial value is 7.50. B. The annual cost is shown by the linear function y = 92.50x. The rate of change is 92.50, and the initial value is 0. C. The annual cost is shown by the linear function y = 7.50x + 85. The rate of change is 7.50, and the initial value is 85. D. The annual cost is shown by the linear function y = x + 85. The rate of change is 1, and the initial value is 85.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Is that it @Summer67894 ?

OpenStudy (anonymous):

Yes...why??

OpenStudy (anonymous):

Let me know if you get any wrong.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

I got three questions wrong but they werent the ones that we did

OpenStudy (anonymous):

Okay, good :)

OpenStudy (anonymous):

But I passed thank you